# TEMPLE: Temporal Preference Learning of Video LLMs via Difficulty Scheduling and Pre-SFT Alignment

### v2

@@ Abstract
Video Large Language Models (Video LLMs) have achieved significant success by leveraging a two-stage paradigm: pretraining on large-scale video-text data for vision-language alignment, followed by supervised fine-tuning (SFT) for task-specific capabilities. However, existing approaches struggle with temporal reasoning due to weak temporal correspondence in the data and reliance on the next-token prediction paradigm during training. To address these limitations, we propose TEMPLE (TEMporal Preference Learning), a systematic framework that enhances Video LLMs’ temporal reasoning capabilities through Direct Preference Optimization (DPO). To facilitate this, we introduce an automated preference data generation pipeline that systematically constructs preference pairs by selecting videos that are rich in temporal information, designing video-specific perturbation strategies, and finally evaluating model responses on clean and perturbed video inputs. Our temporal alignment features two key innovations: curriculum learning which that progressively increases perturbation difficulty to improve model robustness and adaptability; and “Pre-SFT Alignment”, applying preference optimization before instruction tuning to prioritize fine-grained temporal comprehension. Extensive experiments demonstrate that our approach consistently improves Video LLM performance across multiple benchmarks with a relatively small set of self-generated DPO data. We further analyze the transferability of DPO data across architectures and the role of difficulty scheduling in optimization. Our findings highlight our TEMPLE as a scalable and efficient complement to SFT-based methods, paving the way for developing reliable Video LLMs. Code is available at https://github.com/lscpku/TEMPLE.
@@ 1 Introduction
Video LLMs have made significant strides in recent years, achieving remarkable success in various video understanding tasks [9, 22]. These advancements follow the standard multi-modal LLM paradigm: First, the model learns visual knowledge and vision-text alignment through large-scale pretraining on image/video-text pairs [19], and then it acquires instruction following capabilities via supervised fine-tuning (SFT) on curated instruction data [18, 29, 39]. This two-stage approach has enabled impressive improvements in video-language modeling, allowing models to generate coherent, contextually relevant responses conditioned on video input [8, 14, 26].
Despite this progress, current Video LLMs still exhibit poor temporal understanding, primarily due to two fundamental limitations. First, constructing high-quality video datasets is inherently challenging due to the complex visual and temporal structures in videos. Existing datasets often suffer from weak video-text correspondence, visual shortcuts, or insufficient temporal reasoning tasks [12, 21]. Second, both pretraining and SFT rely on next-token prediction using ground-truth text, which does not explicitly enforce dynamic temporal understanding. As a result, models frequently neglect subtle yet significant temporal details, over-rely on partial visual or textual cues, and struggle with temporal relationships between events, leading to inconsistent or flawed responses.
@@
To address these limitations, in this paper, we propose TEMPLE (TEMporal Preference LEarning), a systematic framework that tackles both challenges through: (1) a scalable, automatic dataset curation pipeline that generates high-quality temporal preference data without relying on external LLMs, and (2) a tailored temporal alignment methodology that goes beyond next-token prediction. TEMPLE is built on Direct Preference Optimization (DPO) [24], which optimizes the relative quality of model responses by maximizing the reward margin between preferred and non-preferred outputs. Unlike SFT, which requires high-quality, human-verified responses, DPO focuses on relative comparisons rather than absolute correctness—making it particularly effective for video tasks where perfect annotations are difficult to obtain. Subtle temporal aspects like sequence order and motion speed are critical for comprehensive video understanding yet challenging to annotate manually at scale. Our preference-based framework naturally discourages shortcut learning while enabling efficient scaling through automated annotation, making TEMPLE an ideal methodology for enhancing temporal reasoning in Video LLMs.
@@
We implement TEMPLE through an automatic preference data generation pipeline that establishes the foundation for effective temporal alignment in Video LLMs. Inspired by advances in image-based LLMs [33, 41], we generate preference pairs by evaluating model responses under both clean and corrupted visual inputs. Our pipeline consists of three key stages: (1) curating a diverse set of videos that emphasize complex temporal relationships; (2) introducing targeted temporal perturbations—including frame dropping, sequence reversal, and frame shuffling—that precisely target known temporal reasoning weaknesses of current Video LLMs; and (3) constructing preference pairs by contrasting the model’s outputs for clean versus temporally corrupted videos. We also introduce a difficulty factor that quantifies perturbation magnitude to control the sample difficulties. This systematic approach enables us to efficiently generate high-quality contrastive examples that directly highlight temporal understanding capabilities without requiring expensive human annotation.
@@
Our tailored temporal alignment framework consists of two key technical innovations. First, we implement curriculum learning with progressive difficulty lifting, which gradually increases sample complexity during optimization. This controlled exposure to increasingly challenging temporal contrastive pairs enables more effective learning. Second, we introduce a novel “Pre-SFT Alignment” strategy, which reverses the conventional order of model alignment. Traditional approaches apply DPO after instruction-tuning (SFT) [30], assuming the model has already learned basic capabilities. In contrast, we argue that robust temporal understanding should be established first, before learning to follow diverse instructions. Our Pre-SFT Alignment applies DPO first on our curated dataset, forcing the model to develop strong temporal reasoning and video-text correspondence prior to any instruction tuning. These innovations, coupled with our carefully crafted self-sufficient video filtering and preference generation pipeline, distinguish our work from concurrent works like Tarsier 2 [34] and TPO [15], which rely on external LLMs (e.g., ChatGPT for Tariser 2 and GPT-4o for TPO) for data filtering and apply DPO in a post-SFT manner.
@@
Through extensive evaluation, we demonstrate that TEMPLE consistently improves model performance across multiple benchmarks including Video-MME [8], MLVU [40] and Vinoground [36] with only a small number of self-generated DPO pairs—eliminating dependence on external models or APIs. Our ablation studies validate the effectiveness of both technical contributions: the difficulty curriculum significantly stabilizes training and improves final performance, while the Pre-SFT Alignment strategy significantly outperforms the conventional SFT-then-DPO approach, particularly on tasks requiring fine-grained temporal reasoning. Further analysis reveals valuable insights about DPO data transferability between different model architectures and scales, demonstrating the framework’s flexibility. By introducing this automated, structured preference learning pipeline, we advance the state of video LLMs and pave the way for more robust and temporally-aware multi-modal models.
@@ 2 Method
In this section, we first present a preliminary analysis on the detailed captioning behavior of current Video LLMs, highlighting the need for a more effective alignment strategy. We then introduce our DPO-based approach and propose a three-stage pipeline for automatically generating preference data by comparing model responses on clean and perturbed video inputs. As part of this process, we incorporate a difficulty factor to control the magnitude of noise, enabling structured learning. Finally, we present our novel “Pre-SFT Alignment” paradigm, where we apply DPO using captioning-based preference pairs before SFT, enhancing fine-grained temporal alignment in Video LLMs.
@@ 2.1 Preliminary Analysis on Detailed Captioning
We conduct a preliminary analysis of the detailed captioning behavior of Qwen2-VL-7B [31] by manually annotating events in video keyframes and comparing them to events captured in model-generated captions across 19 videos. The comparison reveals a precision of 74.4% and a recall of merely 35.0%, indicating that the model not only introduces a considerable amount of hallucinated content but also overlooks a substantial portion of events present in the videos (example shown in Figure 1). These findings underscore the critical need for improving video LLMs’ ability to capture and retain temporal information accurately.
2.2 DPO Data Generation Pipeline
Video DPO data consists of four key components: video, instruction, chosen response, and rejected response. Inspired by POVID [41], we generate video DPO data by feeding clean and noised visual inputs into Video LLMs, using the response to the clean input as the chosen response and the response to the corrupted input as the rejected response. To enable preference learning without reliance on external LLMs or APIs, we design a structured pipeline for generating video DPO data, which follows a three-stage process to systematically construct videos, craft instructions, and derive response pairs (Figure 2).
2.2.1 Video Selection
@@
In the first step, we perform scene boundary detection using TransNetV2 [27], which offers more accurate segmentation than tools like PySceneDetect [^1] https://github.com/Breakthrough/PySceneDetect, particularly in handling gradual transitions such as fade-ins and fade-outs. The number of detected scenes serves as an indicator of a video’s temporal complexity, helping us distinguish between dynamic and static content. To refine our selection, we discard monochrome scenes shorter than 0.2 seconds, as these typically correspond to transition frames with little meaningful information. We also filter out videos containing scenes longer than 16 seconds, as such segments often consist of unstructured long shots that are difficult to decompose into discrete events, making them unsuitable for preference learning. Since current segmentation tools struggle to effectively break down these complex videos, we discard them in favor of those with clearer event boundaries.
@@
In the second step, we further refine our selection by filtering out videos that contain frequent scene cuts but little new visual information. This includes cases where multiple shots capture the same scene from different angles or repetitive back-and-forth cuts, such as those commonly found in interview footage. To achieve this, we introduce a similarity grouping stage, which computes the cosine similarity between middle frames of clips using SigLIP [35] and iteratively group clips with similarity above a predefined threshold. If a video contains too few unique groups, it is deemed to have low information density and is removed. Conversely, if a video contains too many unique groups, it is considered excessively complex, presenting challenges for both computational efficiency and model capability. To balance these factors, we retain only videos with 4 to 32 distinct groups, ensuring sufficient diversity while maintaining manageable complexity.
@@
In the final step, we extract keyframes that best represent the visual content of each selected clip. Since we have limited clip length to a maximum of 16 seconds, ensuring that each clip corresponds to a single event, we choose to extract two keyframes per clip, which is generally sufficient to capture its core elements. To improve the clarity of the selected keyframes, we set anchor points at one-third and two-thirds of the clip’s duration and select frames that are both close to these anchors and exhibit high visual clarity. Clarity is assessed using the Laplacian operator [11], which measures sharpness to filter out blurry frames. This final step ensures that the extracted keyframes provide distinct and representative snapshots of the video, further enhancing the quality of our DPO data.
@@
Through this three-step process, we systematically filter videos to maximize temporal diversity, eliminate redundant or overly simplistic content, and ensure that selected clips contain clear, well-defined events. This careful curation allows us to construct a dataset that effectively supports Video LLMs in learning robust temporal understanding through preference optimization.
2.2.2 Perturbation Strategy and Task Design
@@
The design of perturbation strategies and task formulation is critical in our DPO data construction pipeline, as they directly influence the model behaviors we aim to reinforce. Given that video-based QA datasets often suffer from low quality and unreliable annotations, we opt to use detailed captioning as the sole task for preference learning. By focusing on captioning rather than QA, we ensure that preference signals are derived from a more structured and reliable task formulation.
@@
Another common failure mode in Video LLMs is the difficulty in maintaining temporal structure and distinguishing between visually similar but sequentially distinct frames. To address this, we introduce random clip shuffling and clip reversal, where clips within a video are either randomly reordered or flipped in temporal sequence. Unlike clip dropping, this perturbation preserves the total amount of visual information but disrupts its logical arrangement. Consequently, the model produces a rejected response that misinterprets event chronology or blends together unrelated segments, highlighting its struggle with temporal reasoning. By optimizing for preference under this perturbation, we encourage the model to develop stronger temporal comprehension and robustness against structural inconsistencies.
These perturbation strategies serve as structured adversarial challenges, enabling the model to learn more precise and reliable video understanding through preference optimization. By constructing preference pairs that reflect real-world failure cases, we ensure that DPO training directly addresses the fundamental shortcomings of current Video LLMs.
2.2.3 Response Generation
@@
Instead of requesting a single caption for the entire video, we employ a contextualized clip captioning approach [3]. Specifically, at each step, we provide the model with two consecutive clips (corresponding to four frames) and prompt it to generate a caption for the latter clip while considering the contextual information from both. This localized approach ensures that captions remain detailed while still maintaining continuity across the video. This strategy effectively injects contextual dependencies into each caption while preventing information overload.
Once individual clip captions are generated, we aggregate them to form a global video-level caption. To achieve this, we feed all clip captions into the model and prompt it to generate a comprehensive, structured summary of the entire video. This multi-stage process ensures that the final caption preserves both fine-grained event details and high-level narrative coherence.
For response generation in the clean input setting, we directly use the model’s output given the full sequence of clip captions. For perturbed inputs, we modify the clip captions according to our designed perturbations—for instance, dropping, shuffling, or reversing some of the clips. The model is then prompted with the altered captions to generate a corresponding response, which serves as the rejected response in our preference learning pipeline. By structuring response generation in this manner, we ensure that our preference pairs reflect meaningful variations in video comprehension, reinforcing the model’s ability to process complex temporal relationships and mitigate common failure cases.
2.3 Difficulty Scheduling
@@
Given a video with $N$ clips, we regulate difficulty as follows: (i) For clip dropping, we retain only $\lceil N/r\rceil$ clips while discarding the rest, reducing the available visual information. (ii) For clip shuffling and reversal, we partition the clips into $\lceil N/r\rceil$ groups, treating each group as an indivisible unit during perturbation. This ensures that lower values of $r$ introduce finer-grained perturbations while preserving more of the original structure, whereas higher values of $r$ cause more drastic disruptions.
Figure 3 shows an illustration of how the videos are perturbed for different difficulty factors. Intuitively, a smaller $r$ results in weaker perturbations, meaning more information and temporal consistency are preserved, making it harder to differentiate between clean and perturbed inputs. Conversely, a larger $r$ leads to stronger perturbations, making distinctions more apparent and simplifying the preference learning task.
@@
2.4 Pre-SFT Alignment
Figure 4: Training loss and gradient norm during SFT. Pre-SFT alignment leads to lower loss and stable gradients during training.
In the conventional training paradigm, DPO is typically applied after SFT. This ordering ensures that the model first acquires the fundamental ability to perform tasks and follow instructions before refining its alignment with human preferences. However, considering our objective of improving the model’s fine-grained understanding of video content and its temporal structure, we propose an alternative strategy: performing DPO before SFT. Our key insight is that the primary limitation of Video LLMs lies in their inability to establish precise correspondences between visual input and textual descriptions. Without a strong foundation in fine-grained video-text alignment, the model is prone to relying on shortcuts and generating responses that are only loosely connected to the visual content. This issue is particularly pronounced in video-based tasks, where temporal dependencies play a crucial role in determining meaning. By applying DPO before SFT, we enable the model to first internalize the intricate relationships between video frames and textual descriptions before moving on to more general instruction-following training. As shown in Figure 4, we observe that models trained with Pre-SFT Alignment exhibit significantly lower loss values and more stable gradient norms during the subsequent SFT phase. This suggests that early-stage temporal alignment enhances the model’s ability to process visual information effectively, leading to improved convergence and overall training stability. By shifting alignment to an earlier stage in training, our approach ensures that Video LLMs develop a robust foundation in temporal understanding, making them better equipped to handle complex multi-modal reasoning tasks.

### v3

# TEMPLE: Incentivizing Temporal Understanding of Video Large Language Models via Progressive Pre-SFT Alignment

@@ Abstract
Video Large Language Models (Video LLMs) have achieved significant success by adopting the paradigm of large-scale pre-training followed by supervised fine-tuning (SFT). However, existing approaches struggle with temporal reasoning due to weak temporal correspondence in the data and over-reliance on the next-token prediction paradigm, which collectively result in the absence temporal supervision. To address these limitations, we propose TEMPLE (TEMporal Preference LEarning), a systematic framework that enhances temporal reasoning capabilities through Direct Preference Optimization (DPO). To address temporal information scarcity in data, we introduce an automated pipeline for systematically constructing temporality-intensive preference pairs comprising three steps: selecting temporally rich videos, designing video-specific perturbation strategies, and evaluating model responses on clean and perturbed inputs. Complementing this data pipeline, we provide additional supervision signals via preference learning and propose a novel Progressive Pre-SFT Alignment strategy featuring two key innovations: a curriculum learning strategy which progressively increases perturbation difficulty to maximize data efficiency; and applying preference optimization before instruction tuning to incentivize fundamental temporal alignment. Extensive experiments demonstrate that our approach consistently improves Video LLM performance across multiple benchmarks with a relatively small set of self-generated DPO data. Our findings highlight TEMPLE as a scalable and efficient complement to SFT-based methods, paving the way for developing reliable Video LLMs.
Codehttps://github.com/lscpku/TEMPLE
@@ 1 Introduction
Video LLMs have made significant strides in recent years, achieving remarkable success in various video understanding tasks (gemini; gpt4o). These advancements follow the standard multi-modal LLM paradigm: First, the model learns visual knowledge and vision-text alignment through large-scale pretraining on image/video-text pairs (liu2023llava), and then it acquires instruction following capabilities via supervised fine-tuning (SFT) on curated instruction data (liu2023llava15; tong2024cambrian; llava-video-sft). This two-stage approach has enabled impressive improvements in video-language modeling, allowing models to generate coherent, contextually relevant responses conditioned on video input (videomme; li2023mvbench; timechat).
Despite this progress, current Video LLMs still exhibit poor temporal understanding, primarily due to two fundamental limitations. First, constructing high-quality video datasets is inherently challenging due to the complex visual and temporal structures in videos. Existing datasets often suffer from weak video-text correspondence, visual shortcuts, or insufficient temporal reasoning tasks (SingleFrameBias; mangalam2023egoschema; vitatecs). Second, both pretraining and SFT rely on next-token prediction using ground-truth text, which does not explicitly enforce dynamic temporal understanding. As a result, models frequently neglect subtle yet significant temporal details, over-rely on partial visual or textual cues, and struggle with temporal relationships between events, leading to inconsistent or flawed responses.
@@
To address these limitations, we propose TEMPLE (TEMporal Preference LEarning), a systematic framework that tackles both challenges through: (1) a scalable data pipeline that generates high-quality temporal preference data without relying on external LLMs, and (2) a tailored temporal alignment methodology that goes beyond next-token prediction. TEMPLE is built on Direct Preference Optimization (DPO) (rafailov2023direct), which optimizes the relative quality of model responses by maximizing the reward margin between preferred and non-preferred outputs. Unlike SFT, which requires high-quality, human-verified responses, DPO focuses on relative comparisons rather than absolute correctness, making it particularly effective for video tasks where perfect annotations are difficult to obtain. By learning from temporality-oriented preference pairs, our framework naturally discourages shortcut learning while enabling efficient scaling through automated annotation.
@@
We implement TEMPLE through an automatic preference data pipeline that addresses the scarcity of temporal supervision in video data. Inspired by advances in image-based LLMs (yu2024rlaif; povid), we generate preference pairs by evaluating model responses under both clean and corrupted visual inputs. Our pipeline consists of three key stages: (1) curating a diverse set of videos that emphasize complex temporal relationships; (2) designing perturbation strategies that precisely target known temporal reasoning weaknesses of current Video LLMs; and (3) constructing preference pairs by contrasting the model’s outputs for clean versus temporally corrupted videos. We also introduce a difficulty factor that quantifies perturbation magnitude to control the sample difficulties. This systematic approach enables us to efficiently generate high-quality contrastive examples that directly highlight temporal understanding capabilities without requiring expensive human annotation.
@@
To leverage the signals in the generated preference data for more effective temporal learning, we further propose a novel Progessive Pre-SFT Alignment strategy tailored for our temporal preference pairs, which consists of two key technical innovations. First, we implement curriculum learning by gradually increasing sample difficulty during preference optimization. This controlled exposure to increasingly challenging temporal contrastive pairs enables more efficient utilization of the preference data. Second, we introduce a novel “Pre-SFT Alignment” strategy, which reverses the conventional order of model alignment. Traditional approaches apply DPO after instruction-tuning (tunstall2023zephyr), assuming the model has already learned basic capabilities. In contrast, we argue that robust temporal understanding should be established first, before learning to follow diverse instructions. These innovations, coupled with our carefully crafted self-sufficient video filtering and preference generation pipeline, distinguish our work from recent works like Tarsier2 (tarsier2) and TPO (tpo), which rely on proprietary LLMs for data filtering and apply DPO in a post-SFT manner.
@@
Through extensive evaluation, we demonstrate that TEMPLE consistently improves model performance across multiple benchmarks including Video-MME (videomme), MLVU (MLVU) and Vinoground (vinoground) with only a small number of self-generated DPO pairs. Ablation studies validate the effectiveness of our Progressive Pre-SFT Alignment strategy, which outperforms both the SFT baseline and the conventional SFT-then-DPO approach. Further analysis reveals the transferrability of our DPO data across different model architectures and scales, demonstrating the framework’s flexibility. By introducing this automated, structured preference learning pipeline, we advance the state of Video LLMs and pave the way for more robust and temporally-aware multi-modal models.
@@ 2 Method
In this section, we first present a preliminary analysis on the detailed captioning behavior of current Video LLMs, highlighting the need for a more effective alignment strategy. We then introduce our DPO-based approach and propose a three-stage pipeline for automatically generating preference data by comparing model responses on clean and perturbed video inputs. Finally, we present Progressive Pre-SFT Alignment, where we apply DPO using captioning-based preference pairs with increasing difficulty before SFT, enhancing fine-grained temporal alignment in Video LLMs.
@@ 2.1 Preliminary Analysis on Detailed Captioning
We conduct a preliminary analysis of the detailed captioning behavior of Qwen2-VL-7B (Qwen2VL) by manually annotating events in video keyframes and comparing them to events captured in model-generated captions across 19 videos. The comparison reveals a precision of 74.4% and a recall of merely 35.0%, indicating that the model not only introduces a considerable amount of hallucinated content but also overlooks a substantial portion of events present in the videos (example shown in Figure 1). These findings underscore the critical need for improving Video LLMs’ ability to capture and retain temporal information accurately.
2.2 Preference Data Generation Pipeline
Video preference data consists of four key components: video, instruction, chosen response, and rejected response. Inspired by POVID (povid), we generate video DPO data by feeding clean and noised visual inputs into Video LLMs, using the response to the clean input as the chosen response and the response to the corrupted input as the rejected response. To enable preference learning without reliance on external LLMs or APIs, we design a structured pipeline for generating video DPO data, which follows a three-stage process to systematically select videos, craft instructions, and derive response pairs (Figure 2).
Video Selection.
@@ 2.2 Preference Data Generation Pipeline
In the first step, we perform scene boundary detection using TransNetV2 (transnetv2), which offers more accurate segmentation than tools like PySceneDetect [^1] https://github.com/Breakthrough/PySceneDetect, particularly in handling gradual transitions such as fade-ins and fade-outs. The number of detected scenes serves as an indicator of a video’s temporal complexity, helping us distinguish between dynamic and static content. To refine our selection, we discard monochrome scenes shorter than 0.2 seconds, as these typically correspond to transition frames with little meaningful information. We also filter out videos containing scenes longer than 16 seconds, as such segments often consist of unstructured long shots that are difficult to decompose into discrete events, making them unsuitable for preference learning. Since current segmentation tools struggle to effectively break down these complex videos, we discard them in favor of those with clearer event boundaries.
@@
In the second step, we further filter out videos that contain frequent scene cuts but little new visual information. This includes cases where multiple shots of the same scene from different angles or repetitive back-and-forth cuts, such as those commonly found in interview footage. To achieve this, we introduce a similarity grouping stage, which computes the cosine similarity between middle frames of clips using SigLIP (siglip) and iteratively group clips with similarity above a predefined threshold. If a video contains too few unique groups, it is deemed to have low information density and is removed. Conversely, if a video contains too many unique groups, it is considered excessively complex, presenting challenges for both computational efficiency and model capability. To balance these factors, we retain only videos with 4 to 32 distinct groups, ensuring sufficient diversity while maintaining manageable complexity.
@@
In the final step, we extract keyframes that best represent the visual content of each selected clip. Since we have limited clip length to a maximum of 16 seconds, ensuring that each clip corresponds to a single event, we choose to extract two keyframes per clip, which is generally sufficient to capture its core elements. To improve the clarity of the selected keyframes, we set anchor points at one-third and two-thirds of the clip’s duration and select frames that are both close to these anchors and exhibit high visual clarity. Clarity is assessed using the Laplacian operator (machine_vision), which measures sharpness to filter out blurry frames. This final step ensures that the extracted keyframes provide distinct and representative snapshots of the video, further enhancing the quality of our DPO data.
@@
Through this three-step process, we systematically filter videos to maximize temporal diversity, eliminate redundant or overly simplistic content, and ensure that selected clips contain clear, well-defined events.
Perturbation Strategy and Task Design.
@@
The design of perturbation strategies and task formulation is critical in our DPO data construction pipeline, as they directly influence the model behaviors we aim to reinforce. Given that video-based QA datasets often suffer from low quality and unreliable annotations, we opt to use detailed captioning as the sole task for preference learning.
@@
Another common failure mode is the difficulty in maintaining temporal structure and distinguishing between visually similar but sequentially distinct frames. To address this, we introduce random clip shuffling and clip reversal, where clips within a video are randomly reordered or flipped temporally. This perturbation preserves the total amount of visual information but disrupts its logical arrangement. Consequently, the model produces a rejected response that misinterprets event chronology or blends together unrelated segments. By optimizing for preference under this perturbation, we encourage the model to develop stronger temporal comprehension and robustness against structural inconsistencies.
These perturbation strategies serve as adversarial challenges, enabling the model to learn more precise and reliable temporal understanding through preference optimization.
Response Generation.
@@
Instead of requesting a single caption for the entire video, we employ a contextualized clip captioning approach (chen2024sharegpt4video). Specifically, at each step, we provide the model with two consecutive clips (corresponding to four frames) and prompt it to generate a caption for the latter while considering the contextual information from both. This localized approach ensures that captions remain detailed while maintaining continuity across the video.
Once individual clip captions are generated, we aggregate them to form a global video-level caption by feeding all clip captions into the model and prompt it to generate a comprehensive summary of the entire video. This multi-stage process ensures that the final caption preserves both fine-grained event details and high-level narrative coherence.
For response generation in the clean input setting, we directly use the model’s output given the full sequence of clip captions. For perturbed inputs, we modify the clip captions according to our designed perturbations. The model is then prompted with the altered captions to generate a corresponding response, which serves as the rejected response in our preference learning pipeline. By structuring response generation in this manner, we ensure that our preference pairs reflect meaningful variations in video comprehension, reinforcing the model’s ability to process complex temporal relationships and mitigate common failure cases.
Figure 4: Training loss and gradient norm during SFT. Pre-SFT Alignment leads to lower loss and more stable gradients during training.
2.3 Progressive Pre-SFT Alignment
Progressive Diffuculty Scheduling.
@@ 2.3 Progressive Pre-SFT Alignment
Given a video with NN clips, we regulate difficulty as follows: (i) For clip dropping, we retain only $\lceil N/r\rceil$ clips while discarding the rest, reducing the available visual information. (ii) For clip shuffling and reversal, we partition the clips into $\lceil N/r\rceil$ groups, treating each group as an indivisible unit during perturbation. This ensures that lower values of $r$ introduce finer-grained perturbations while preserving more of the original structure, whereas higher values of $r$ cause more drastic disruptions.
Figure 3 illustrates how the videos are perturbed for different difficulty factors. Intuitively, a smaller $r$ results in weaker perturbations, meaning more information and temporal consistency are preserved, making it harder to differentiate between clean and perturbed inputs. Conversely, a larger $r$ leads to stronger perturbations, making distinctions more apparent and simplifying the preference learning task.
@@
Pre-SFT Alignment.
In the conventional training paradigm, DPO is typically applied after SFT. This ordering ensures that the model first acquires the fundamental ability to perform tasks and follow instructions before refining its alignment with human preferences. However, considering our objective of improving the model’s fine-grained understanding of video content and its temporal structure, we propose an alternative strategy: performing DPO before SFT. Our key insight is that the primary limitation of Video LLMs lies in their inability to establish precise correspondences between visual input and textual descriptions. Without a strong foundation in fine-grained video-text alignment, the model is prone to relying on static shortcuts and generating responses that are only loosely connected to the true dynamics of the visual content. By applying DPO before SFT, we enable the model to first internalize the intricate relationships between video frames and textual descriptions before moving on to more general instruction-following training. As shown in Figure 4, we observe that models trained with Pre-SFT Alignment exhibit significantly lower loss values and more stable gradient norms during the subsequent SFT phase. This suggests that early-stage temporal alignment enhances the model’s ability to process visual information effectively, leading to improved convergence and overall training stability. By shifting alignment to an earlier stage in training, our approach ensures that Video LLMs develop a robust foundation in temporal understanding, making them better equipped to handle complex multi-modal reasoning tasks.
Model # Params VideoMME (TP/TR) MLVU Vinoground_txt TempCompass_mc Average
Gemini-1.5-pro - 75.0 - 35.8 63.9 -
GPT-4o - 77.2 64.6 54.0 80.0 67.7
VideoLLaMA2 72B 62.4 61.2 36.2 - -
TimeMarker 7B 57.3 63.9 - 65.8 -
Video-T3 7B 55.0 58.1 31.6 61.2 51.5
LongVA 7B 52.6 56.3 21.2 56.1 46.6
LLaMA-3-VILA1.5 8B 45.9 22.4 - 56.4 -
LLaVA-NeXT-Video 34B 60.2 - 23.0 68.7 -
LLaVA-OneVision 7B 58.2 64.7 41.6 64.8 57.3
MiniCPM-V-2.6 8B 60.9 - 32.6 63.0 -
Video-LLaVA 7B 39.9 47.3 24.8 45.6 39.4
Qwen2-VL-2B 2B 55.2 (54.5/34.5) 57.7 26.4 64.6 51.0
   + SFT 2B 54.6 (65.5/32.8) 60.0 27.4 65.0 51.8
   + TEMPLE (Ours) 2B 55.1 (63.6/35.0) 60.2 28.0 64.6 52.0
Qwen2-VL-7B 7B 60.4 (70.9/42.4) 64.5 34.0 69.6 57.1
   + SFT 7B 60.7 (70.9/44.1) 64.6 34.6 70.8 57.7
   + TEMPLE (Ours) 7B 61.0 (74.5/45.2) 65.8 37.2 71.4 58.9
Table 1: Benchmark accuracy of our approach on VideoMME, MLVU, Vinoground (Text), and TempCompass (Multi-Choice). TP and TR denotes Temporal Perception and Temporal Reasoning, respectively, two subtasks of VideoMME.